OpenStudy (anonymous):

Terry has completed 15 homework problems, and Susan has completed 9. Terry is completing 10 problems per hour, and Susan is completing 12 per hour. After how many hours will Terry and susan have completed the same number of problems?

OpenStudy (anonymous):

3hours 4hours 5hours 10hours

OpenStudy (anonymous):

5 hours I believe, let me double-check my work.

OpenStudy (anonymous):

Made a mistake actually, it's 4 hours. @Youth

OpenStudy (anonymous):

(1 hour) Terry = 15 Susan = 9 ( 2 hours) Terry = 25 Susan = 21 (3 hours) Terry = 35 Susan = 33 (4 hours) Terry = 45 Susan = 45